To calculate a colleague's graduation, the m workers in an office agreed to contribute equally to a catered lunch that costs a total of y dollars. If p of the coworkers fail to contribute, which of the following represents the additional amount, in dollars that each of the remaining coworkers must contribute to pay for the lunch?

A: y/m
B: y/(m-p)
C: (py)/(m-p)
D: (y(m-p))/m
E: (py)/(m(m-p))

Answer by ptaylor(2198)


If all the workers contributed, each would contribute y/m dollars. But since P of the coworkers failed to contribute, then each of the remaining coworkers contributed y/(m-p) dollars. So we have:
(y/(m-p))-y/m)=the additional amount that each would have contribute
Simplifying we get: (multiply numerator and denominator of each term by m(m-p)
(ym-y(m-p))/m(m-p)=
ym-ym+yp/m(m-p)=
yp/m(m-p)

Hope this helps---ptaylor